OpenStudy (anonymous):

how can I solve this? 1/x + 5/25 = 3/5 Solve for x

OpenStudy (anonymous):

You could subtract 5/25 from both sides and then use cross multiplication..

OpenStudy (lilg132):

1/x = 3/5 - 5/25 1/x = 2/5 5/x = 2 5 = 2x 5/2 = 2x / 2 2.5 = x Dont forget to click best response !